OpenStudy (anonymous):

Can some one help me I need to learn how to solve this I know I need to use y = mx + b A doctor measured a baby’s length at 21 inches when the baby was one month old. At the baby’s six-month doctor’s appointment, the baby’s length was 26 inches. Find the baby’s length when the baby is one year old if his length increases at a constant rate.

OpenStudy (anonymous):

31 inches i think? this suppose to be a A.P term

OpenStudy (campbell_st):

ok... if you want to use the straight line y = mx + b then you need to find the slope using the given data points (1, 21) and (6, 26). are you able yo calculate the slope for the 2 points above..?

OpenStudy (campbell_st):

@Zheng... why not let the asker... answer the questions...

OpenStudy (anonymous):

So do point slope formula with what cambell provided...do you know how to do that?

OpenStudy (anonymous):

31 is the answer. am I rgt?

OpenStudy (anonymous):

No your answer is not a number

OpenStudy (anonymous):

Its going to be in terms of x

OpenStudy (campbell_st):

@rambo... I have no idea

OpenStudy (campbell_st):

actually its incorrect @rambo

OpenStudy (anonymous):

Do (26-21)/(6-1) to do point slope formula, I basically just gave you the answer

OpenStudy (anonymous):

@ilikephysics2 you are wrong. yo can get number.

OpenStudy (anonymous):

No you can't its in y = mx+b

OpenStudy (campbell_st):

lol..@ilikephysics2 your are busting to answer this...

OpenStudy (anonymous):

i guess you can make it a number, or leave it in terms of x

OpenStudy (anonymous):

Cambell you have no idea what you are doing do you?

OpenStudy (campbell_st):

you are asked to use the linear relationship y = mx + b to answer the question @ilikephysics2 , not just get the get the equation

OpenStudy (anonymous):

so use it in that formula, thats all she needs

OpenStudy (campbell_st):

actually I do know how to solve this question

OpenStudy (anonymous):

okay guys -.- thanks lol

OpenStudy (anonymous):

btw i was trying to help her do the formula not just giving her the answer

OpenStudy (campbell_st):

why not wait as see if she knows what to do...rather that saying... this is it... but its all just maths

OpenStudy (anonymous):

let me explain you. y=mx+b y=lenght m=month x,b = constant at the time of birth, y=21 m=1 21=x+b at 6th mont, y=26 m=6 26=6x+b we have x+b=21 6x+b=26 by solvinf this 2, we get, x=1 b=20 now at one year, y=? m=12 x=1 b=20 y=12*1 + 20 y=32

OpenStudy (anonymous):

hope every1 got the answer.